OpenStudy (lgbasallote):

another one i got wrong... \[\large \int_5^{\infty} \frac{2dx}{x\sqrt{x^2 - 25}}\]

OpenStudy (lgbasallote):

my answer was \(\large \frac{\pi}{5}\) and got 6 points out of 12 =_=

OpenStudy (lgbasallote):

@myininaya ...need a consult

myininaya (myininaya):

\[\int\limits_{}^{}\frac{2 dx}{x \sqrt{x^2-25}}\] Did you have a problem integrating?

OpenStudy (lgbasallote):

uhmm you can judge for yourself... \[2\lim_{t \rightarrow \infty} \int_5^t \frac{dx}{x\sqrt{x^2 - 25}}\]

OpenStudy (lgbasallote):

\[2[\frac{1}{5} \sec^{-1} (\frac{x}{5})|_5^t]\]

OpenStudy (lgbasallote):

\[2[\frac{1}{5} \sec^{-1} (\infty)] - 2[\frac{1}{5} \sec^{-1} (1)]\]

OpenStudy (lgbasallote):

\[2[\frac{1}{5} (\frac{\pi}{2})] - 2[\frac{1}{5} (0)]\]

myininaya (myininaya):

\[\int\limits \frac{2 dx}{x \sqrt{x^2-25}}\]  I labeled it this way so that i would have what is inside that square rooty :) \[\sec(\theta)=\frac{x}{5} => \sec(\theta) \tan(\theta) d \theta=\frac{1}{5} dx\] So we have \[2 \int\limits_{}^{}\frac{5 \sec(\theta) \tan(\theta) d \theta}{5 \sec(\theta) \sqrt{(5\sec(\theta))^2-25}}\] \[\frac{2}{5}\int\limits_{}^{}\frac{\sec(\theta) \tan(\theta) d \theta}{\sec(\theta)\sqrt{\sec^2(\theta)-1}}\] \[\frac{2}{5}\int\limits_{}^{}1 d \theta=\frac{2}{5}\theta+C=\frac{2}{5}\sec^{-1}(\frac{x}{5})+C\] Ok I approved of your integrating part

myininaya (myininaya):

lol

OpenStudy (lgbasallote):

lol :p so where did i go wrong

myininaya (myininaya):

\[\int\limits_{5}^{52} f(x) dx+\int\limits_{52}^{\infty}f(x) dx\] I just split them up at some random constant between 5 and infinity lol \[[\lim_{a \rightarrow 5^+}\frac{2}{5}\sec^{-1}(\frac{x}{5})]_a^{52}+[\lim_{b \rightarrow \infty}\frac{2}{5} \sec^{-1}(\frac{x}{5})]_{52}^{b}\]

myininaya (myininaya):

sorry it keeps going to that aw snap page so annoying

OpenStudy (lgbasallote):

it is isnt it

myininaya (myininaya):

\[\lim_{a \rightarrow 5^+}\frac{2}{5}(\sec^{-1}(\frac{52}{5})-\sec^{-1}(\frac{a}{5}))+\lim_{b \rightarrow \infty} \frac{2}{5}(\sec^{-1}(\frac{b}{5})-\sec^{-1}(\frac{52}{5}))\] Now do both parts converge?

OpenStudy (lgbasallote):

by converge means definite value right?

OpenStudy (lgbasallote):

seems like it

myininaya (myininaya):

a finite value yes and yes you are right

myininaya (myininaya):

lol

myininaya (myininaya):

\[\frac{2}{5}(\sec^{-1}(\frac{52}{5})-0)+\frac{2}{5}(\frac{\pi}{2}-\sec^{-1}(\frac{52}{5}))\] \[\frac{2}{5} \cdot \frac{\pi}{2}\]

myininaya (myininaya):

pi/5 is right

OpenStudy (lgbasallote):

then why did i get 6 out of 12 points =_=

myininaya (myininaya):

maybe she/he didn't like the way that you didn't split it up

OpenStudy (lgbasallote):

oh lol

myininaya (myininaya):

there was 2 different discont.\ infinite situation going on there

OpenStudy (lgbasallote):

finkledipper -_-

myininaya (myininaya):

so you need to do 2 separate thingys lol

OpenStudy (lgbasallote):

well thanks :D

myininaya (myininaya):

you're welcome

myininaya (myininaya):

by the way if you haven't noticed i don't like using the formulas i like do it the long way every time lol

OpenStudy (lgbasallote):

haha yeah i noticed :p i did that too but too much latex and drawings so i just wrote it directly hehe